How many permutations can be made out of the letters of the word 'BUSINESS'? How many of these begin with B and end with N ?
Answer by stanbon(75887)

How many permutations can be made out of the letters of the word 'BUSINESS'?
Ans: 8!/(3!) = 8*7*6*5*4 = 6720
-----------------------------
How many of these begin with B and end with N ?
-----
Ans: 1*(6!/2!)*1 = 360
-------------------------------
Cheers,
Stan H.